A physics book slides off a horizontal tabletop with a speed of 3.5 m/s. It strikes the floor in 0.350 s. Ignore air resistance.
Use g = 10 m/s2.
Find:

a. The horizontal and vertical components of the velocity when the book leaves the table.

b. The height of the tabletop above the floor (how far does it travel vertically?).

c. The horizontal distance from the edge of the table to the point where the book strikes the floor (how far does it travel horizontally?).

d. The horizontal and vertical components of the book’s velocity just before the book reaches the floor.
Answer #1

Answer0 35 sec ره) Vertical component= Vy = y + gt Given Speed = 3.5m/s Time= Horizontal component= 3.5m ls Vertical component=0 As